OpenStudy (princessashley):

What is the average rate of change of f(x), represented by the table of values, over the interval [-3, 4]? x f(x) -6 27 -3 6 -1 2 0 3 1 6 4 27

OpenStudy (latinc):

is this for physics

OpenStudy (zzr0ck3r):

The average rate of change of the function \(f\) over \([a,b]\) is given by \[\dfrac{f(a)-f(b)}{a-b}\]

OpenStudy (princessashley):

okay how do i find fa and fb? do i just test the numbers from the chart?

OpenStudy (sshayer):

f(a)=f(-3)=6 f(b)=f(4)=27 b-a=4-(-3)=4+3=7 f(b)=27 f(a)=6 average rate of change=\[\frac{ f(b)-f(a) }{ b-a }\]=?

OpenStudy (triciaal):

Substitute the values in the formula and compute

OpenStudy (triciaal):

the average rate of change is like finding the slope over the specified period of change

OpenStudy (triciaal):

can you close now or do you still have a question?